# Active Learning of Model Discrepancy with Bayesian Experimental Design

### v1

@@ 4 Discussions
This work adopts a Bayesian framework rooted in information theory while aligning more closely with traditional experimental design due to the reliance on measured data. Specifically, we conduct preliminary experiments with several proposed designs, evaluate the information gained from each, and use this information to optimize for an improved design. While the likelihood and posterior updates rely on an inaccurate model, the results presented in Section 3 demonstrate that this approach yields highly informative designs with respect to effectively enabling model correction despite discrepancies. The optimization problem can be practically addressed by proposing several candidate designs and selecting or interpolating the most informative one. Alternatively, it can be solved using either gradient-based methods or derivative-free optimization techniques. It is worth noting that the numerical results of Section 3 only employ the measurement data of the optimal design point for model error correction, instead of the measurement data on a trajectory of points that lead to the optimal design point. If the whole trajectory of measurement data is used for model error correction, the performance of the subsequent posterior distribution estimation for the physics-based model parameters could be slightly worse, which seems to be counter-intuitive and is mainly due to the direct contribution of the optimal design point in the subsequent posterior distribution estimation. This accounts for the main reason that having a better match of measurement data on the optimal design point is more favorable than an overall good match of measurement data on the whole trajectory of searching the optimal design point.
@@
In the numerical results of Section 3, we mainly focused on the scenario that allows some experimental measurements during the iterative stages of sequential experimental design. We have also tested the scenario that solely depends on the numerical model simulations, and the comparisons indicate that both predicted and measured data are effective within our framework. The use of simulation data yields less satisfactory performances and demands further technical refinement. The reason for this performance gap is that the EIG of a given design can differ significantly from the actual information gain it provides. When the model is correct, this difference is primarily determined by the accuracy of the prior information. However, in the presence of model discrepancy, errors in the likelihood function further amplify this difference, leading to a scenario where the design identified based on EIG may yield an incorrect information gain. The biased likelihood function is utilized twice in the computation of basic version EIG: (i) to calculate the predicted distribution of measurements, and (ii) to update the posterior distribution after obtaining a specific prediction. By using real measurements, the first step can be bypassed, thereby mitigating the impact of model inaccuracies. Consequently, the computed information gain represents a more reliable metric of information provided by the data. To demonstrate the strength of the hybrid correction framework, we focus on using measurement data in our design process and leave the scenario that solely depends on numerical model simulations to approximate EIG for future work.
@@ 5 Conclusion
In this work, we propose an efficient hybrid approach to actively learning the model discrepancy of a digital twin based on the data from the sequential BED. We first validate the performance of the proposed method in a classical numerical example governed by a convection-diffusion equation, for which full BED is still feasible. The results confirm that the proposed method can effectively calibrate the low-dimensional parametric model error based on the optimal designs from the BED. The proposed method is then further studied in the same numerical example with a high-dimensional structural model error, for which full BED is not practical anymore. We observe that most of the optimal designs from the BED focus on the low-dimensional physics-based parameters can still be informative for calibrating the model discrepancy, while some of those designs could gradually lead to a larger model discrepancy. Therefore, we propose an ensemble-based approximation of information gain to assess the data informativeness and to enhance the active learning model discrepancy. The results show that the proposed method is efficient and robust to active learning of the high-dimensional model discrepancy. With the promising results of the hybrid approach demonstrated in this work, some future directions are summarized below:
@@ Acknowledgments
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Appendix A Algorithm
A detailed algorithm for sequential BED with model discrepancy correction and ensemble-based indicator has been summarized in Algorithm 1.
Algorithm 1 Active Learning of Model Discrepancy with BED
for $i=1,2,...,N$ do $\triangleright$ Iterating stage
     $\mathbf{d}\leftarrow\operatorname*{arg\,max}_{\mathbf{d}\in\mathcal{D}}\mathbb{E}[U(\bm{\theta}_{\mathcal{G}},\mathbf{y},\mathbf{d};\bm{\theta}_{\textrm{NN}})]$triangleright $\bm{\theta}_{\textrm{NN}}$ is fixed
     $\mathbf{y}\leftarrow\mathbf{u}(\mathbf{d})+\bm{\epsilon}$triangleright Measurement at optimal design
     $\bm{\theta}_{\mathcal{G}}\leftarrow\bm{\theta}_{\mathcal{G}}|\mathbf{y}$triangleright Bayesian update
     $\bm{\theta}_{\mathcal{G}}^{*}\leftarrow\operatorname*{arg\,max}_{\bm{\theta}_{\mathcal{G}}}p(\bm{\theta}_{\mathcal{G}};\bm{\theta}_{\text{NN}})$triangleright MAP
         $\bm{\theta}_{\text{NN}}\leftarrow\operatorname*{arg\,max}_{\bm{\theta}_{\text{NN}}}L(\bm{\theta}_{\text{NN}};\bm{\theta}_{\mathcal{G}}^{*},\mathbf{y},\mathbf{d})$triangleright $\bm{\theta}_{\mathcal{G}}^{*}$ is fixed
     end if
end for
Appendix B Theoretical global maximum
@@
With $\bm{\theta}_{\mathcal{G}}^{*}$ approximating $\bm{\theta}_{\mathcal{G}}^{\dagger}$, a result progressively provided by Bayesian optimal experimental design (OED)[26, 66], the above alignment equation can represent the equivalence between Eq. (1) and Eq. (2). The universal approximation property of neural networks guarantees the existence of a solution, which corresponds to the optimized $\bm{\theta}_{\text{NN}}^{*}$.
@@
    $\bm{\theta}_{\mathcal{G}}^{\dagger}$ can be approximated by $\bm{\theta}^{*}_{\mathcal{G}}$ from the posterior of $\Theta_{\text{OED}}$ by BED;
@@
Appendix C Detailed derivation of the Jacobian matrix
@@
=$=\left.\left\{\nabla^{2}\eta(\mathbf{z})-v(t)\cdot\nabla\eta(\mathbf{z})\right\}_{i}\right|_{h=0}$
=$=\left\{\left[\nabla^{2}-v(t)\cdot\nabla\right]\eta(\mathbf{z})\right\}_{i}$

### v2

@@ 3.2 Hybrid Approach on Structural Errors
3.3 Hybrid Approach on Ill‑Posed Problems
In this section, we present a more challenging numerical example that aims at illustrating the performance of our method in an ill‑posed setting, i.e., the solution of the physics-based parameters and the neural-network-based structural error is not unique. Extending from the structural error case presented in Section 3.2, we further assume incorrect prior knowledge in the linearly increasing convection velocity $v(t)$. Specifically, the accurate velocity on $x$-axis is known, whereas the growth rate $k_{y}$ in $v_{y}=k_{y}t$ is known as a wrong value. Intuitively, a stronger convective component along one axis deflects the passive scalar plume toward the orthogonal direction, producing an asymmetrically skewed trajectory in the scalar field. This setting increases the nonlinearity of the inverse problem, since the convection coefficient $k_{y}$ enters the advective operator $v_{y}\partial_{y}=(k_{y}t)\partial_{y}$, making the map from $k_{y}$ to the observed field values inherently nonlinear rather than a simple scalar scaling of the data.
The source location $\{\theta_{x},\theta_{y}\}$ remains unknown, and the previously assumed incorrect source form is retained under network correction $\bm{\theta}_{\text{NN}}$. Following the workflow in Section 3.2, this problem can be solved by treating $\{\theta_{x},\theta_{y},k_{y}\}$ as physical parameters and introducing a network with $\bm{\theta}_{\text{NN}}$ to capture the structural error. To study an ill-posed scenario, only the parameters $\{\theta_{x},\theta_{y}\}$ are modeled as random variables characterized by probability distributions and are utilized for identifying optimal design, while the parameters $\{k_{y},\bm{\theta}_{\text{NN}}\}$ are updated jointly via gradient-based methods. The ill-posedness arises from the flexibility of the neural network, i.e., the network is sufficiently expressive to match the observed data even when $k_{y}$ takes an incorrect value. Consequently, many pairs of $\{k_{y},\bm{\theta}_{\text{NN}}\}$ can yield simulation results that match the observed data. It is worth noting that the ill-posedness discussed here is different from that in Section 3.2, i.e., the trained neural network can compensate or even completely replace the role of existing physics-based terms in the modeled system, thus making $\theta_{\mathcal{G}}$ essentially non-identifiable. In this section, we focus on studying the possible ill-posedness in the optimization stage, which is an internal non-identifiability that emerges during the joint optimization of certain physical parameters (e.g., $k_{y}$) with the network parameters ($\theta_{\text{NN}}$). This challenging scenario is also significantly exacerbated by the limited amount of training data at every optimization stage. We first demonstrate the ill-posedness experimentally by fixing $k_{y}$ at an incorrect value and only updating $\bm{\theta}_{\text{NN}}$. Subsequently, we illustrate that the inability to precisely identify the true value of $k_{y}$ does not harm the inference of $\{\theta_{x},\theta_{y}\}$.
For the detailed setup, the true source parameters $\{\theta_{x},\theta_{y}\}$ is set as $\{0.25,0.35\}$ with a uniform prior distribution on $0,1]^{2}$. The true value of the growth rate $k_{y}$ is $50$ while the initially assumed incorrect value is $70$. The structural error in the source term, the neural network architecture, and other conditions remain consistent with those in Section 3.2. Instead of employing a single measurement location, we adopt a centered $3\times 3$ sensor grid consisting of the original point and its eight immediate neighbors at intervals of 0.06. This increases the total amount of information in the data, thus alleviating the intrinsic challenge of jointly learning $k_{y}$ and the structural error. This setup retains physical interpretability, as sensor arrays in small regions or a rapidly movable sensor (with measurement times shorter than system evolution times) are both common in the experimental design community.
We first consider the case in which $k_{y}$ is fixed at an incorrect value and $\theta_{\text{NN}}$ is updated. Figure 10 presents the inference results of $\{\theta_{x},\theta_{y}\}$ using different models. Without the correction of $\theta_{\text{NN}}$, as shown in Figure 10(a), the model with biased convection speed and incorrect source form yields a biased and multi-modal posterior distribution. However, with an updating network $\bm{\theta}_{\text{NN}}$ and fixed biased $k_{y}$, the highlighted areas in Figure 10(b) progressively converge toward the true source location, demonstrating that our framework accurately identifies the source position. We attribute this improvement to the neural network for compensating the incorrect convection term as well. To verify this argument, we compute the relative errors of the simulated solution fields and the PDE right-hand side (RHS) as defined in Eq. (23), excluding the diffusion term, at various update stages before and after correction. The results with fixed $k_{y}$ in Table 2 show that the data-driven correction consistently reduces the RHS error, even with an incorrect $k_{y}$ value, thereby enhancing the accuracy of the simulated fields in the vicinity of the data. Optimizing $k_{y}$ allows for directly adjusting the incorrect physical convection characteristics of the modeled system, leading to more significant error reduction. It is expected that the neural network designed to model a spatially decaying source correction term is inherently limited in its ability to globally compensate for an inaccurate convection velocity, compared to the direct adjustment of $k_{y}$.
Table 2: Relative errors of the simulated solution fields and the PDE right-hand-side (RHS) without diffusion at different update stages. “No” refers to the model before applying data-driven correction at the current stage, while “Yes” refers to the corrected model after incorporating data at the same stage. The RHS error corresponds to the right-hand side of the PDE, excluding the diffusion term, as the diffusion term is same between model and system. All errors are computed over a small region slightly larger than the observation area.
Stage Index 2 3 4 5
Model Correction No Yes No Yes No Yes No Yes
Fix $k_{y}$
RE of simulated field 0.4951 0.1544 0.3575 0.0687 0.2552 0.1534 0.7752 0.2742
RE of RHS* 0.6378 0.6248 0.5069 0.4514 0.6224 0.5895 0.6971 0.6968
Update $k_{y}$
RE of simulated field 0.4951 0.1003 0.2343 0.1069 0.7091 0.1037 0.2461 0.1159
RE of RHS* 0.6378 0.5664 0.6479 0.6122 0.4831 0.3555 0.4225 0.4413
(a) With $k_{y}$ and $\theta_{\text{NN}}$ frozen
(b) With $k_{y}$ frozen
(c) Update $k_{y}$
Figure 10: Posterior distributions of inferred source location $\{\theta_{x},\theta_{y}\}$ with (a) both $k_{y}$ and $\theta_{\text{NN}}$ fixed, (b) $k_{y}$ fixed and $\bm{\theta}_{\text{NN}}$ updated , and (c) $k_{y}$ and $\theta_{\text{NN}}$ jointly updated. In panel (a), no model correction is applied; only one design point is selected per stage using BED, resulting in a single observation per plot. In panel (b) and (c), larger orange dots indicate design points used for BED, while smaller semi-transparent orange dots denote additional observations used for model correction.
We further investigate the detailed results of the PDE RHS (excluding the diffusion term) and the solution field across successive update stages. In Fig. 11(a), the results indicate that the modeled right-hand side is gradually matching the true system within the collected-data region. Consequently, in Fig. 11(b), the solutions of the modeled system (second row) gradually converge to the true field (first row), and the overall residual is steadily reduced, demonstrating that the solution field is effectively corrected. This confirms that the neural-network-corrected RHS remains close to the ground truth, yielding accurate posterior inferences overall, although $k_{y}$ does not correctly converge to the value taken by the true system.
These results confirm that our approach effectively aligns the modeled system with the true one. This neural-network-corrected modeled system thus ensures an accurate and robust posterior estimation of the physics-based parameters treated as random variables and updated via Bayesian methods. It should be noted that treating some of the physics-based parameters deterministically and optimizing them jointly with neural network parameters can cause the neural network to compensate the incorrect values taken by those physics-based parameters, impairing the identifiability of those parameters and leading to a challenging ill-posed problem.
(a) Results for the RHS of the PDE
(b) Results for the solution field of the PDE
Figure 11: Comparison between the true data, modeled approximation, and residuals across five measurement stages for RHS and solution field of the PDE: (a) Results for the RHS of the PDE, and (b) Results for the solution field of the PDE. Each panel consists of three rows: the first row shows the true data, the second row displays the modeled approximation, and the third row presents the residuals (difference between the true and modeled values). The contour plots are zoomed into the region $0.1,0.5]^{2}$, which corresponds to the primary data collection area for model calibration. Within each panel, identical colors represent the same numerical values throughout the entire figure. For modeled results, stage 1 shows the initial model before any update, while subsequent stages present the updated model at each corresponding stage. In the first stage, model correction is skipped, and only the design point for the physical parameter is shown.
@@ 4 Discussions
This work adopts a Bayesian framework rooted in information theory while aligning more closely with traditional experimental design due to the reliance on measured data. Specifically, we conduct preliminary experiments with several proposed designs, evaluate the information gained from each, and use this information to optimize for an improved design. While the likelihood and posterior updates rely on an inaccurate model, the results presented in Section 3 demonstrate that this approach yields highly informative designs with respect to effectively enabling model correction despite discrepancies. The optimization problem can be practically addressed by proposing several candidate designs and selecting or interpolating the most informative one. Alternatively, it can be solved using either gradient-based methods or derivative-free optimization techniques. It is worth noting that the numerical results of Section 3 only employ the measurement data of the optimal design point for model error correction, instead of the measurement data on a trajectory of points that lead to the optimal design point. Admittedly, more data along the whole optimization trajectory tends to provide more information gain, while the calibration of the model error can possibly be worse by using the trajectory of data points, and thus the criterion based on information gain to assess the data informativeness on correcting the model error needs to build on a fixed amount of data points, e.g., a single data point in this work. The reason for only adopting the data at the optimal design point for a given stage is mainly due to the direct contribution of the optimal design point in the posterior distribution estimation of the subsequent stage. Therefore, a better match of measurement data on the optimal design point is more favorable than an overall good match of measurement data on the whole optimization trajectory, for which the model accuracy on the optimal design point could be compensated a bit to ensure the overall agreement with the whole trajectory data.
A key consideration in hybrid models that combine physics-based components with neural-network-based correction terms is the ill-posedness, where the network may compensate or even replace existing physics-based components, thereby obscuring the actual meaning of those physics-based parameters being estimated. In certain cases, such as the numerical example studied in Section 3.2, such ill-posedness can be mitigated by imposing desired constraints on the trained neural network, e.g., by enforcing the prescribed center-decay pattern with assigned center and restricting the network to spatial variation only, while the time-varying convective term is handled entirely by the physics-based model. More broadly, general cases may require other strategies like promoting orthogonality or post-training decomposition. An example of post-training decomposition is a re-attribution step to find an adjusted estimate $\bm{\theta}_{\mathcal{G}}^{\text{adj}}$ by requiring the physical model $\mathcal{G}(\mathbf{u};\bm{\theta}_{\mathcal{G}})$ to maximally explain the learned RHS $H(\mathbf{u})=\mathcal{G}(\mathbf{u};\bm{\theta}_{\mathcal{G}}^{*})+\text{NN}(\mathbf{u};\bm{\theta}_{\text{NN}})$, via solving an additional optimization problem $\bm{\theta}_{\mathcal{G}}^{\text{adj}}=\arg\max_{\bm{\theta}_{\mathcal{G}}}\|\mathcal{G}(\mathbf{u};\bm{\theta}_{\mathcal{G}})-H(\mathbf{u})\|^{2}_{2}$. The $\bm{\theta}_{\mathcal{G}}^{\text{adj}}$ is thus an updated estimation of the physics-based parameters where all aspects of the trained RHS explainable by the physics-based model $\mathcal{G}$ are attributed to it. Although this post-training decomposition is not investigated in this work, it leads to a meaningful future direction for the joint learning of physics-based parameters and structural errors.
@@
In the numerical results of Section 3, we mainly focused on the scenario that allows some experimental measurements during the iterative stages of sequential experimental design. We have also tested the scenario that solely depends on the numerical model simulations, and the comparisons indicate that both predicted and measured data are effective within our framework. The use of simulation data yields less satisfactory performance and demands further technical refinement. The reason for this performance gap is that the EIG of a given design can differ significantly from the actual information gain it provides. When the model is correct, this difference is primarily determined by the accuracy of the prior information. However, in the presence of model discrepancy, errors in the likelihood function further amplify this difference, leading to a scenario where the design identified based on EIG may yield incorrect information gain. The biased likelihood function is utilized twice in the computation of basic version EIG: (i) to calculate the predicted distribution of measurements, and (ii) to update the posterior distribution after obtaining a specific prediction. By using real measurements, the first step can be bypassed, thereby mitigating the impact of model inaccuracies. Consequently, the computed information gain represents a more reliable metric of information provided by the data. To demonstrate the strength of the hybrid correction framework, we focus on using measurement data in our design process and leave the scenario that solely depends on numerical model simulations to approximate EIG for future work.
In our framework, we partition the parameters into two groups and handle each group with a distinct update mechanism, which raises the question of how to select a representative point estimate from a probabilistic posterior for use in the deterministic gradient updates. We address this by performing multiple gradient-based optimizations to find distinct local maxima, a strategy that allows us to more effectively synthesize information from the posterior landscape. We then use the mean of these identified peaks as the averaged point estimate for the physics-based parameters. We have demonstrated that this strategy is both effective and robust when the true posterior is a unimodal distribution, or at least dominated by a leading mode. If the true posterior of the physics-based parameters is apparently multimodal, averaging of samples may yield an estimate that lies between peaks and degrades network performance at each mode. In such cases, it is potentially advantageous to train different neural networks at each representative mode and then weight their contributions by distance or likelihood, which is a meaningful extension for future study.
@@ 5 Conclusion
In this work, we propose an efficient hybrid approach to actively learning the model discrepancy of a digital twin based on the data from the sequential BED. We first validate the performance of the proposed method in a classical numerical example governed by a convection-diffusion equation, for which full BED is still feasible. The results confirm that the proposed method can effectively calibrate the low-dimensional parametric model error based on the optimal designs from the BED. The proposed method is then further studied in the same numerical example with a high-dimensional structural model error, for which full BED is not practical anymore. It is worth noting that this numerical example could suffer from the ill-posedness issue, which stems from jointly learning the physics-based model parameters and the neural-network-based correction term. Such an ill-posedness issue is mitigated in this work by restricting the network to learn only the spatial discrepancy, while delegating all time-dependent terms to the physics-based model. We observe that most of the optimal designs from the BED that focus on the low-dimensional physics-based parameters can still be informative for calibrating the model discrepancy, while some of those designs could gradually lead to a larger model discrepancy. Therefore, we propose an ensemble-based approximation of information gain to assess the data informativeness and to enhance the active learning model discrepancy. The results show that the proposed method is efficient and robust to active learning of the high-dimensional model discrepancy. With the promising results of the hybrid approach demonstrated in this work, some future directions are summarized below:
@@ Acknowledgments
The codes and data that support the findings of this study are available in the link: https://github.com/AIMS-Madison/BED-Model-Discrepancy.
Appendix A Theoretical global maximum
@@
With $\bm{\theta}_{\mathcal{G}}^{*}$ approximating $\bm{\theta}_{\mathcal{G}}^{\dagger}$, a result progressively provided by Bayesian OED[27, 70], the above alignment equation can represent the equivalence between Eq. (1) and Eq. (2). The universal approximation property of neural networks guarantees the existence of a solution, which corresponds to the optimized $\bm{\theta}_{\text{NN}}^{*}$.
@@
    $\bm{\theta}_{\mathcal{G}}^{\dagger}$ can be approximated by $\bm{\theta}^{*}_{\mathcal{G}}$ from the its posterior by BED;
@@
Appendix B Detailed derivation of the Jacobian matrix
@@
$=\left.\left\{\nabla^{2}\eta(\mathbf{z})-v(t)\cdot\nabla\eta(\mathbf{z})\right\}_{i}\right|_{h=0}$
$=\left\{\left[\nabla^{2}-v(t)\cdot\nabla\right]\eta(\mathbf{z})\right\}_{i}$
@@
Appendix C Additional results of an acoustic inverse problem
The wave equation:
$\frac{\partial^{2}u(\mathbf{r},t)}{\partial t^{2}}-c^{2}\nabla^{2}u(\mathbf{r},t)=0,\quad\mathbf{r}\in\Omega$ (33)
describes how a physical quantity propagates through space and time. For the steady-state or single-frequency case, we assume that the solution has a harmonic time dependence. By substituting an ansatz of the form
$u(\mathbf{r},t)=U(\mathbf{r})e^{i\omega t},$ (34)
into the wave equation, the time-dependent part factors out. This process leads to Helmholtz equation for the spatial component:
$\nabla^{2}U(\mathbf{r})+k^{2}U(\mathbf{r})=0,$ (35)
where the wave number $k=\omega/c=2\pi/\lambda$, with $c$ represents the phase speed and $\lambda$ represents the wave length. The complex acoustic pressure amplitude function $U(\mathbf{r})$ is complex-valued, which represents both the amplitude and phase information at each point $\mathbf{r}$. The Helmholtz equation is central to understanding the spatial behavior of waves under monochromatic conditions, making it invaluable in fields such as acoustics, electromagnetics, and quantum mechanics.
We consider the problem of a point source acoustic wave propagating in a two-dimensional inhomogeneous medium [73, 74]:
$-\nabla^{2}U(x,y)-e^{2m(x,y)}k^{2}U(x,y)$ $=f,\quad\{x,y\}\in\Omega,$ (36)
$f$ $=A\delta(x-x_{0},y-y_{0}),$
$\nabla u\cdot\mathbf{n}$ $=0~\text{on}~\Gamma_{\text{top}},$
PML boundary condition $~\text{on }~\partial\Omega/\Gamma_{\text{top}}.$
where $e^{2m(x,y)}$ represents the spatially varying medium property, such as the local reflection coefficient or refractive index contrast. $f$ represents a time-harmonic point source of amplitude $A$, located at $\{x_{0},y_{0}\}$ and modeled by Dirac delta function $\delta$. The computational domain $\Omega$ is a square region, subject to a reflective (Neumann) boundary condition on the top boundary $\Gamma_{\text{top}}$ and surrounded by perfectly matched layers (PML) on the remaining three boundaries to simulate wave absorption. Due to the inhomogeneity, the medium’s physical properties vary with position, which leads to complex propagation behavior. Increasing $m(x,y)$ reduces the local wave speed and increases the effective refractive index, resulting in stronger reflection and denser wavefronts with more rapid spatial oscillations. For example, Fig. 12(a) shows the time-domain wave field snapshot of the case with $m(x,y)=2/3x$, a high wave number $k=100$, and a source located in $\{x_{0},y_{0}\}=\{0.75,1.25\}$, where the wavefronts to become progressively denser in regions with larger $x$-coordinate. Also, the reflection can be clearly seen near the top boundary.
(a) Example wave field at time unit 0
(b) True magnitude field of the complex amplitude
Figure 12: Field results of the acoustic system. Panel (a) is the time-domain snapshot of wave field at time unit 0 with k=100. Panel (b) is the magnitude of the complex amplitude $U$ with k=5, which is true field in the inverse problem. The white dots denote the potential measure places.
Problem setup: Assume the unknown physical parameter is the source amplitude $A$ with true value $1.5$, and a model discrepancy lies in the medium properties. The true form of $m^{\dagger}(x,y)$ in system is given by
$m^{\dagger}(x,y)=\frac{2}{3}x,$ (37)
and the modeled $m(x,y)$ is
$m(x,y)=\frac{2}{3}x^{2}.$ (38)
A neural network (denoted by NN) is utilized to capture the discrepancy between the modeled and true medium properties: $\text{NN}\rightarrow\frac{2}{3}(x-x^{2})$. The network takes the spatial coordinate $x$ as input and outputs a scalar medium parameter $m$ at spatial location $\{x,y\}$: $\text{NN}:x\mapsto m$. The architecture is a fully connected feed-forward network featuring a single hidden layer with five nodes, which results in a total of 16 parameters $\bm{\theta}_{\text{NN}}\subseteq\mathbb{R}^{16}.$
The general goal of this case is to seek a design $\mathbf{d}$ as where to measure $U$ and use the measurement $\mathbf{y}=U(\mathbf{d})+\epsilon,~\epsilon\sim\mathcal{N}(0,\sigma^{2})$ to infer the source amplitude $A$. A brief procedure for each sequential experiment is described by
  • 1.
    Find optimal $\mathbf{d}$ for $A$ by BED method.
  • 2.
    Observe $\mathbf{y}$ and update a posterior for $A$: $p(A|\mathbf{y},\mathbf{d})$ with a MAP estimation $A^{*}$.
  • 3.
    Optimize the network parameter $\bm{\theta}_{\text{NN}}$ by maximizing likelihood $p(\mathbf{y}|\mathbf{d},A^{*},\bm{\theta}_{\text{NN}})$.
  • 4.
    Use the updated model to re-update a posterior $p(A|\mathbf{y},\mathbf{d})$.
The posterior of each experiment will be utilized as the prior for the next experiment. And the updated model is also passed into the next experiment.
Numerical Setting and Solver: The computational domian $\Omega$ is defined as $0,1.5]^{2}$, discretized uniformly into $64\times 64$ cells. The wave number is set to $k=5.0$ and the source is located at $\{x_{0},y_{0}\}=\{0.75,1.25\}$. The potential measure places are constrained to the set $\{\frac{3}{128}+\frac{3}{64}i,\frac{189}{128}\}_{i=0}^{31}$ as shown in Fig. 12(b). This setup follows the cases in [73, 74], but adopts a scaled-down version. The scaling is intentional: while the referenced works focus on tractable estimation of high-dimensional expected information gain, our emphasis lies in learning and correcting model discrepancy. Despite the simplified scaling, the simulated wave fields closely resemble those presented in the original works under the same setting. The prior distribution of $A$ is assumed to be uniform over $[0,3]$, discretized into 151 equally spaced points. A finite element method solver is developed entirely in JAX for its automatic differentiation capabilities. The demonstration example will be available in our GitHub repository.
Results: We begin by applying the standard BED to find design and inference source amplitude using an uncorrected model. The results, presented in the first row of Fig. 13 show that the posterior distributions remain biased due to the persistent impact of model discrepancy on parameter inference. Next, we implement our proposed approach, in which the model is updated using the collected data, and the posterior distribution is subsequently re-evaluated. As illustrated in the second row of Fig. 13, the posterior distributions progressively converge toward the true parameter values and become more concentrated, indicating that the model correction effectively enhances parameter inference. We further check the performance of the updated network. The comparison of true $m^{\dagger}$, initial wrong $m$, and the corrected $m+\text{NN}$ is shown in Fig. 14. Within the data region, the network-corrected model $m+\text{NN}$ aligns much more closely with the true model $m^{\dagger}$, thereby demonstrating the informativeness of the data optimized for $A$ and the effectiveness of the correction approach. This improved model accuracy further enhances design optimization and parameter inference, accounting for the fine results in Fig. 13.
Figure 13: Posterior distribution of source amplitude AA
Figure 14: Comparison of Medium Property Fields: True $m^{\dagger}(x)$, Initial $m(x)$ and Corrected $m(x)+\text{NN}(x|\bm{\theta}_{\text{NN}})$. The data points represents the $x$-coordinates corresponding to the optimal design along with the associated true values $m^{\dagger}$. However, it is important to note that the data used in model correction are not these $\{x,m^{\dagger}\}$ pairs; rather, they are the complex amplitude measurements obtained at the design coordinates.